# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v2

@@ Appendix D Proofs of Section 5
for $\mu_{0,y_{i}}$-a.e. $x\in\mathbb{R}^{d}$. $\Box$
@@ Appendix E Proofs of Section 6
[…]
and thus the claim. $\Box$
In order to proof Proposition 12 we need some auxiliary results, in particular the following lemma which is a variant of [6, Proposition 7.1.3].
Lemma 17.
@@
Proof.
Let $f:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $(y,x)\mapsto(\sqrt{\beta}y,x)$. Then for $\mu_{1},\mu_{2}\in P_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we have that $W_{2,d_{\beta}}(\mu_{1},\mu_{2})=W_{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ since there is a bijection of couplings
@@
which implies that optimal couplings are mapped to optimal couplings. Since also $f_{\sharp}\mu_{n}\rightharpoonup f_{\sharp}\mu,f_{\sharp}\nu_{n}\rightharpoonup f_{\sharp}\nu$ we can use [6, Proposition 7.1.3] to guarantee the existence of a subsequence of optimal plans $\tilde{\alpha}_{n_{k}}$ for $W_{2}(f_{\sharp}\mu_{n_{k}},f_{\sharp}\nu_{n_{k}})$ such that $\tilde{\alpha}_{n_{k}}\rightharpoonup\tilde{\alpha}$ for an optimal plan $\tilde{\alpha}$ for $W_{2}(f_{\sharp}\mu,f_{\sharp}\nu)$. Thus by (107) there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n},\nu_{n})$ such that $\alpha_{n_{k}}\rightharpoonup\alpha$ for an optimal plan $\alpha$ for $W_{2,d_{\beta}}(\mu,\nu)$. ∎
Remark 18.
A useful well known observation is the following. Let $\alpha_{n}\in\Gamma(\mu,\nu),n\in\mathbb{N}$ where $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$. Assume that we have weak convergence $\alpha_{n}\rightharpoonup\alpha$ for some $\alpha\in\mathcal{P}(\mathbb{R}^{2l})$. Then $\alpha\in\Gamma(\mu,\nu)$ and for any measurable function $f:\mathbb{R}^{2l}\to\mathbb{R}$ such that $|f(x_{1},x_{2})|\leq\|x_{1}\|^{2}+\|x_{2}\|^{2}$ we have that $\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha_{n}\to\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha$. The latter claim follows from [6, Remark 5.2.3] which implies that $\|x_{1}\|^{2}+\|x_{2}\|^{2}$ is uniformly integrable w.r. to $\{\alpha_{n},n\in\mathbb{N}\}$ since it has fixed marginals $\mu,\nu$ with finite second moments. Then [6, Lemma 5.1.7], which states that $\lim_{n\to\infty}\int f\mathrm{d}\alpha_{n}\to\int f\mathrm{d}\alpha$ for uniformly integrable $f$ and weak converging $\alpha_{n}\rightharpoonup\alpha$, implies the latter claim. It is immediate that $\alpha\in\Gamma(\mu,\nu)$ since $\Gamma(\mu,\nu)$ is compact and thus closed in the weak topology by [6, Remark 5.2.3].
Note that [31, Proposition 3.11] states the following proposition only under some regularity conditions on $\mu,\nu$ which ensures uniqueness of optimal plans. But this is not needed in their proof if one is only interested in the existence of a optimal limit point. For the convenience of the reader we include a proof adapted to our situation but claim no originality.
Proposition 19.
@@
Proof.
By [6, Remark 5.2.3] the set $\Gamma(\mu,\nu)$ is compact with respect to the weak convergence topology and thus also the closure of $\{\alpha^{\beta_{k}}:k\in\mathbb{N}\}$ is compact. Consequently accumulation points exist. Let $\alpha\in\Gamma(\mu,\nu)$ be an accumulation point and by abuse of notation let $\alpha^{\beta_{k}}\rightharpoonup\alpha\in\Gamma(\mu,\nu)$. Then by Remark 18 also
@@
Since the limes on the left is 00 by Proposition 10 we know that $\pi^{1,3}_{\sharp}\alpha$ is only supported on the diagonal. Thus for any $\pi^{1,3}_{\sharp}\alpha$ measurable bounded function $f:\mathbb{R}^{2d}\to\mathbb{R}$ we have that $f=f\circ\Delta\circ\pi^{1}$ for $\pi^{1,3}_{\sharp}\alpha$ a.e. $(y_{1},y_{2})\in\mathbb{R}^{2d}$ and hence
@@
Hence $\alpha$ is an optimal plan for $W_{2,Y}(\mu,\nu)$ and thus the claim. ∎
Proof of Proposition 12. By [13, Theorem 8.3.2] we know that weak convergence of probability measures is metrizable and we denote by $d_{weak}$ a metric on $\mathcal{P}\left(\mathbb{R}^{2d}\times\mathbb{R}^{2m}\right)$ that metrizes weak convergence. Let $\alpha_{n}$ be a an optimal plan for $\mu_{n},\nu_{n}$ and $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$. Then by [6, Proposition 7.1.3] and Lemma 17, there exists a subsequence of $\alpha_{n}$ optimal for $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$ converging weakly to an optimal plan $\alpha^{\beta_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu,\nu)$. Thus we can find a sequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu_{n_{k}},\nu_{n_{k}})$ such that $d_{weak}(\alpha^{\beta_{k}},\alpha_{n_{k}})<\frac{1}{k}$. We know by [31, Proposition 3.11] resp. Proposition 19 that $\alpha_{\beta_{k}}\to\alpha$ with respect to $d_{weak}$ for an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$. Thus, for $\epsilon>0$, there exists a $k$ such that $\frac{1}{k}+d_{weak}(\alpha^{\beta_{k}},\alpha)<\epsilon$ and we obtain
@@
which proves the claim. $\Box$
@@ Appendix F Implementation Details
We use a setup similar to [55], using the time dependent U-Net architecture from [46] which are trained using Adam [36]. As in [55] we clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. The differences are we use a constant learning rate of 2e-4, 256 model channels and no dropout. We train using 50k target samples for 500 epochs using a batch size of 500 for the minibatch OT couplings and a batch size of 100 for training the networks. We set the same random seed during training to be able to compare runs for different sources of couplings. The conditional coupling plans are calculated using the Python Optimal Transport package [22]. For inference simulate the corresponding ODEs using the torchiffeq [17] package. To evaluate our results, we use the Fréchet inception distance (FID) [30] [^2] We use the implementation from https://github.com/mseitzer/pytorch-fid.. We compute the distance on 50k training samples, for which we generate 50k samples given the same labels as the training samples.

### v3

@@ Appendix D Proofs of Section 5
for $\mu_{0,y_{i}}$-a.e. $x\in\mathbb{R}^{d}$. $\blacksquare$
Proof of Proposition 9
In this paragraph we give a precise statement of Proposition 9 as well as its proof. Recall that convex domain $\Omega\subset\mathbb{R}^{n}$ is called uniformely convex, if for every $\varepsilon>0$ there exists a $\delta=\delta(\varepsilon)>0$ such that for any two points $x,y\in\Omega$ satisfying $\|x-y\|\geq\varepsilon$, the midpoint $m=\frac{x+y}{2}$ fulfills $\mathrm{dist}\big{(}m,\partial\Omega\big{)}\geq\delta$. Here $\partial\Omega$ denotes the boundary of $\Omega$. Furthermore we say that a function $f:\Omega\to B$ for $\Omega\subset\mathbb{R}^{n}$ open and $B$ a Banach space, is a $C^{1}$ map if it is continously Frechet differentiable.
Assumption 1. We say that a measure $P_{Y}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ fulfills Assumption 1, if there exists a uniformly convex bounded open $C^{2}$ subdomain $\Omega_{Y}\subseteq\mathbb{R}^{d}$ such that $P_{Y}(\Omega_{Y})=1$ and it admits a density $p_{Y}\in C^{2}(\Omega_{Y})$ such that there exists $0<\delta<\epsilon$ such that $\delta\leq p_{Y}(y)\leq\epsilon$ for all $y\in\Omega_{Y}$.
Assumption 2. A measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ is said to fulfill Assumption 2 if $\mu=P_{Y}\times\mu^{Z}$ for $\mu^{Z}\in\mathcal{P}_{2}(\mathbb{R}^{m})$ a measure such that there exists a uniformly convex bounded open $C^{2}$ subdomain $\Omega_{Z}\subseteq\mathbb{R}^{m}$ such that $\mu_{Z}(\Omega_{Z})=1$ and it admits a density $p_{Z}\in C^{2}(\Omega_{Z})$ such that there exists $0<\delta<\epsilon$ such that $\delta\leq p_{Z}(z)\leq\epsilon$ for all $z\in\Omega_{Z}$.
Assumption 3 A measure $\mu\in P_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ is said to fulfill Assumption 3 if there exists a disintegration $\mu=\mu^{y}\times_{y}P_{Y}(y)$ such that there exists a uniformly convex bounded open $C^{2}$ subdomain $\Omega\subseteq\mathbb{R}^{m}$ such that $\mu^{y}(\Omega)=1$ and it admits a density $p^{y}\in C^{2}(\Omega)$ such that there exists $0<\delta<\epsilon$ such that $\delta\leq p^{y}(x)\leq\epsilon$ for all $x\in\Omega$.
Proposition 17
Let $P_{Y}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ satisfy Assumption 1, $\mu_{0}=P_{Y}\times\mu^{Z}_{0}$ satisfy Assumption 2 and let $\mu_{1}=\mu_{1}^{y}\times_{y}P_{Y}$ satisfy Assumption 3 with density $p_{1}^{y}$ of $\mu_{1}^{y}$ on $\Omega$. Assume further that the map $y\mapsto p_{1}^{y}:\Omega_{Y}\to C^{2}(\Omega)$ be a $C^{1}$ map. Then there exists a $W_{2,Y}$-optimal transport map $T:(y,x)\mapsto(y,T_{y}(x))$ i.e. $\alpha=(\operatorname{Id},T)_{\sharp}\mu_{0}\in\Gamma_{o,Y}(\mu_{0},\mu_{1})$ where $T_{y}$ is the optimal transport map for $\mu_{0}^{Z}$ and $\mu_{1}^{y}$. Let $\mu_{t}=(e_{t})_{\sharp}\alpha$ with associated vector field $v_{t}\in L^{2}_{\mu_{t}}$, where $(v_{t})_{j}=0$ for all $j\leq d$. Then there is a representative of $v_{t}$ such that the flow equation
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}=v_{t}(\phi_{t});\quad\phi_{0}(y,x)=(y,x)$ (107)
admits a global solution and $\mu_{t}=(\phi_{t})_{\sharp}\mu_{0}$. Furthermore, we have
$v_{t}(\phi_{t}(y,x))=T(y,x)-(y,x)=\left(0,T_{y}(x)-x\right)$ (108)
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$.
Proof  We will first construct a vector field which describes the inverse curve starting in $\mu_{1}$ and ending in $\mu_{0}$ first. Let $T^{y}$ be the $C^{1}$ Monge map between $\mu_{1}^{y}$ and $\mu_{0}^{Z}$, which exists and is unique by the Caffarelli’s regularity Theorem, see (Villani, 2009, Theorem 12.50, ii)). In order to use the latter theorem we need assumptions 2 and 3. Note that we have that $T_{t}^{y}(x)\coloneqq(1-t)x+tT^{y}(x)$ is a invertible $C^{1}$ map from $\Omega_{1}$ onto its image by the proof of Proposition 16 ii). Using assumptions 1-3 the assumption that $y\mapsto p_{1}^{y}$ is $C^{1}$, by (González-Sanz and Sheng, 2024, Corollary 1.2) we have that $T(y,x)\coloneqq(y,T^{y}(x))$ is continuous. Thus we can conclude that $T$ is measurable and hence it is a Monge map with respect to $W_{2,Y}$ for $\mu_{1}$ and $\mu_{0}$. More precisely we have that $(\operatorname{Id},T)_{\sharp}\mu_{1}\in\Gamma_{o,Y}(\mu_{1},\mu_{0})$ which follows e.g. from (10).
Claim: For $t\in[0,1]$ the map $T_{t}(y,x)=(y,T_{t}^{y}(x))$ as map $T_{t}:\Omega_{1}\times\Omega_{Y}\to\mathbb{R}^{d+m}$ is continuous and injective and its image, denoted by $O_{t}\subset\mathbb{R}^{d+m}$ is a Borel set. The continuity follows from the continuity of $T$ and the injectivity from the injectivity of the individual $T_{t}^{y}$. The image of $\Omega_{1}\times\Omega_{Y}$ is Borel measurable as image of an open set under a injective continuous map.
Claim: Let $\nu_{t}=T_{t,\sharp}\mu_{1}$. Then $\nu_{t}=\mu_{1-t}$ and $O_{t}\subseteq\operatorname{supp}(\nu_{t})$ as well as $\nu_{t}(O_{t})=1$. These claims follow from straigthforward computations
Claim: Let $O$ be the image of $(\operatorname{Id},T_{t}):[0,1]\times\Omega_{1}\times\Omega_{Y}\to[0,1]\times\mathbb{R}^{d+m}$. Then $O$ is Borel measurable and we can define a Borel measurable map $T_{t}^{-1}:O\to\mathbb{R}^{d+m}$ which we can view as element in $L^{2}(\int\nu_{t}\mathrm{d}t,\mathbb{R}^{d+m})$. This follows since $(\operatorname{Id},T_{t})$ is continuous and injective and thus has Borel measurable image. Furthermore it is injective which is why we can invert it on its image.
Claim: The map $u_{t}(y,x)\coloneqq(T-\operatorname{Id})\circ T_{t}^{-1}$ is well defined as function in $L^{2}(\int\nu_{t}\mathrm{d}t)$. This follows from above.
Claim: For $\phi_{t}=T_{t}$ we have that $\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}(y,x)=u_{t}(\phi_{t}(y,x))$ and
$u_{t}\nu_{t}=e_{t,\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))(\operatorname{Id},T)_{\sharp}\mu_{1})\right).$
Both claims can be verified from straightforward computations.
Claim: Then $v_{t}\coloneqq-u_{1-t}$ and $T^{-1}$ fulfill the claim. Note that one can easily see that $T^{-1}$ is a Monge map for $\mu_{0}$ and $\mu_{1}$. Furthermore $T^{-1}(y,x)=(y,(T^{y})^{-1})$ and $(T^{y})^{-1}$ is a Monge map between $\mu_{0}^{Z}$ and $\mu_{1}^{y}$. By definition we have $\mu_{t}=(T^{-1})_{t,\sharp}\mu_{0}$ for $(T^{-1})_{t}(y,x)=(1-t)(y,x)+tT^{-1}(y,x)=(y,(1-t)x+t(T^{y})^{-1}(x))$. It is easy to see that $(T^{-1})_{t}=T_{1-t}\circ T^{-1}$ and in turn we know that its image is a Borel set with unit mass under $\mu_{t}=\nu_{1-t}$ making $v_{t}\in L^{2}(\mu_{t})$ well defined. Furthermore $\frac{\mathrm{d}}{\mathrm{d}t}(T^{-1})_{t}=T^{-1}-\operatorname{Id}$. Computing
$v_{t}\circ(T^{-1})_{t}=-u_{1-t}\circ(T^{-1})_{t}=-(T-\operatorname{Id})\circ T_{1-t}^{-1}\left(\circ T_{1-t}\circ T^{-1}\right)=T^{-1}-\operatorname{Id}$ (109)
we can conclude the claim.
@@ Appendix E Proofs of Section 6
$\geq\beta\int_{A^{2}}\|y_{1}-y_{2}\|^{p}\,d\pi^{1,3}_{\sharp}\alpha$
and thus the claim. $\blacksquare$
In order to proof Proposition 12 we need some auxiliary results, in particular the following lemma which is a variant of (Ambrosio et al., 2005, Proposition 7.1.3).
Lemma 18
@@
Proof  Let $f:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $(y,x)\mapsto(\sqrt{\beta}y,x)$. Then for $\mu_{1},\mu_{2}\in P_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we have that $W_{2,d_{\beta}}(\mu_{1},\mu_{2})=W_{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ since there is a bijection of couplings
@@
which implies that optimal couplings are mapped to optimal couplings. Since also $f_{\sharp}\mu_{n}\rightharpoonup f_{\sharp}\mu,f_{\sharp}\nu_{n}\rightharpoonup f_{\sharp}\nu$ we can use (Ambrosio et al., 2005, Proposition 7.1.3) to guarantee the existence of a subsequence of optimal plans $\tilde{\alpha}_{n_{k}}$ for $W_{2}(f_{\sharp}\mu_{n_{k}},f_{\sharp}\nu_{n_{k}})$ such that $\tilde{\alpha}_{n_{k}}\rightharpoonup\tilde{\alpha}$ for an optimal plan $\tilde{\alpha}$ for $W_{2}(f_{\sharp}\mu,f_{\sharp}\nu)$. Thus by (110) there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n},\nu_{n})$ such that $\alpha_{n_{k}}\rightharpoonup\alpha$ for an optimal plan $\alpha$ for $W_{2,d_{\beta}}(\mu,\nu)$.
Remark 19
A useful well known observation is the following. Let $\alpha_{n}\in\Gamma(\mu,\nu),n\in\mathbb{N}$ where $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$. Assume that we have weak convergence $\alpha_{n}\rightharpoonup\alpha$ for some $\alpha\in\mathcal{P}(\mathbb{R}^{2l})$. Then $\alpha\in\Gamma(\mu,\nu)$ and for any measurable function $f:\mathbb{R}^{2l}\to\mathbb{R}$ such that $|f(x_{1},x_{2})|\leq\|x_{1}\|^{2}+\|x_{2}\|^{2}$ we have that $\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha_{n}\to\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha$. The latter claim follows from (Ambrosio et al., 2005, Remark 5.2.3) which implies that $\|x_{1}\|^{2}+\|x_{2}\|^{2}$ is uniformly integrable w.r. to $\{\alpha_{n},n\in\mathbb{N}\}$ since it has fixed marginals $\mu,\nu$ with finite second moments. Then (Ambrosio et al., 2005, Lemma 5.1.7), which states that $\lim_{n\to\infty}\int f\mathrm{d}\alpha_{n}\to\int f\mathrm{d}\alpha$ for uniformly integrable $f$ and weak converging $\alpha_{n}\rightharpoonup\alpha$, implies the latter claim. It is immediate that $\alpha\in\Gamma(\mu,\nu)$ since $\Gamma(\mu,\nu)$ is compact and thus closed in the weak topology by (Ambrosio et al., 2005, Remark 5.2.3).
Note that (Hosseini et al., 2024, Proposition 3.11) states the following proposition only under some regularity conditions on $\mu,\nu$ which ensures uniqueness of optimal plans. But this is not needed in their proof if one is only interested in the existence of a optimal limit point. For the convenience of the reader we include a proof adapted to our situation but claim no originality.
Proposition 20
@@
Proof  By (Ambrosio et al., 2005, Remark 5.2.3) the set $\Gamma(\mu,\nu)$ is compact with respect to the weak convergence topology and thus also the closure of $\{\alpha^{\beta_{k}}:k\in\mathbb{N}\}$ is compact. Consequently accumulation points exist. Let $\alpha\in\Gamma(\mu,\nu)$ be an accumulation point and by abuse of notation let $\alpha^{\beta_{k}}\rightharpoonup\alpha\in\Gamma(\mu,\nu)$. Then by Remark 19 also
@@
Since the limes on the left is 0 by Proposition 10 we know that $\pi^{1,3}_{\sharp}\alpha$ is only supported on the diagonal. Thus for any $\pi^{1,3}_{\sharp}\alpha$ measurable bounded function $f:\mathbb{R}^{2d}\to\mathbb{R}$ we have that $f=f\circ\Delta\circ\pi^{1}$ for $\pi^{1,3}_{\sharp}\alpha$ a.e. $(y_{1},y_{2})\in\mathbb{R}^{2d}$ and hence
@@
Hence $\alpha$ is an optimal plan for $W_{2,Y}(\mu,\nu)$ and thus the claim.
Proof of Proposition 12. By (Bogachev and Ruas, 2007, Theorem 8.3.2) we know that weak convergence of probability measures is metrizable and we denote by $d_{weak}$ a metric on $\mathcal{P}\left(\mathbb{R}^{2d}\times\mathbb{R}^{2m}\right)$ that metrizes weak convergence. Let $\alpha_{n}$ be a an optimal plan for $\mu_{n},\nu_{n}$ and $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$. Then by (Ambrosio et al., 2005, Proposition 7.1.3) and Lemma 18, there exists a subsequence of $\alpha_{n}$ optimal for $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$ converging weakly to an optimal plan $\alpha^{\beta_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu,\nu)$. Thus we can find a sequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu_{n_{k}},\nu_{n_{k}})$ such that $d_{weak}(\alpha^{\beta_{k}},\alpha_{n_{k}})<\frac{1}{k}$. We know by (Hosseini et al., 2024, Proposition 3.11) resp. Proposition 20 that $\alpha_{\beta_{k}}\to\alpha$ with respect to $d_{weak}$ for an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$. Thus, for $\epsilon>0$, there exists a $k$ such that $\frac{1}{k}+d_{weak}(\alpha^{\beta_{k}},\alpha)<\epsilon$ and we obtain
@@
which proves the claim. $\blacksquare$
@@ Appendix F Implementation Details
We use a setup similar to (Tong et al., 2023), using the time dependent U-Net architecture from (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021) which are trained using Adam (Kingma and Ba, 2015). As in (Tong et al., 2023) we clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. The differences are we use a constant learning rate of 2e-4, 256 model channels and no dropout. We train using 50k target samples for 500 epochs using a batch size of 500 for the minibatch OT couplings and a batch size of 100 for training the networks. We set the same random seed during training to be able to compare runs for different sources of couplings. The conditional coupling plans are calculated using the Python Optimal Transport package (Flamary et al., 2021). For inference simulate the corresponding ODEs using the torchiffeq (Chen, 2018) package. To evaluate our results, we use the Fréchet inception distance (FID) (Heusel et al., 2017) [^2] We use the implementation from https://github.com/mseitzer/pytorch-fid.. We compute the distance on 50k training samples, for which we generate 50k samples given the same labels as the training samples.